OpenStudy (help108):

MEDAL AND FAN!!!! What is the value of a in the equation a = 2 + 3a + 8?

OpenStudy (anonymous):

Combine like terms so 2 + 8 is 10

OpenStudy (photon336):

what do you think the first step would be? a = 2+3a+8

OpenStudy (help108):

uhhh 13a?

OpenStudy (anonymous):

No

OpenStudy (photon336):

You would get a = 3a+10

OpenStudy (help108):

ok

OpenStudy (photon336):

The next step is to get a by itself how would you do this?

OpenStudy (help108):

OHHHH

OpenStudy (help108):

10 minus 3a? or just move the a

OpenStudy (anonymous):

Let's solve your equation step-by-step. a=2+3a+8 Step 1: Simplify both sides of the equation. a=2+3a+8 Simplify: (Show steps) a=3a+10 Step 2: Subtract 3a from both sides. a−3a=3a+10−3a −2a=10 Step 3: Divide both sides by -2. −2a−2=10−2 a=−5 Answer: a=−5

OpenStudy (photon336):

so you would subtract the 3a from both sides

OpenStudy (anonymous):

U get it

OpenStudy (photon336):

a-3a = 10 now what?

OpenStudy (help108):

@clanger555 just told me how to do it @Photon336

OpenStudy (help108):

clanger just told me how to do it photon

OpenStudy (anonymous):

:)

OpenStudy (help108):

is he right @Photon336

OpenStudy (help108):

@Photon336

OpenStudy (help108):

is he right?

OpenStudy (anonymous):

I think I am he gave me medal and to check replace the a with the answer

OpenStudy (killuaxzoldyck):

a=-5 Is the correct answer. :)

OpenStudy (help108):

thanks

OpenStudy (anonymous):

Np